1.4 Given the constraints (10) A+B + C <= 24, B +C >=8 and A >= 0, B >= 0, C>=  0. Maximize 24-A-B - C A: amount of time spent on school work B: amount of time spent on fun C: amount of time spent on pay work
1
Expert's answer
2020-05-10T12:33:56-0400

Maximize "Z=24-A-B-C"

subject to,

"A+B+C \le 24\\
B+C\ge8\\
A,B,C\ge0"


for simplex calculation,


Maximize "Z=24-A-B-C+0S_1+0S_2-MA_1"

subject to,

"A+B+C+S_1 = 24\\
B+C-S_2+A_1=8\\
A,B,C,S_1,S_2,A_1\ge0"


iteration 1


iteration 2

answer is ,

"\bold{A=0,B=8,C=0}\\
Z_{max}=24-0-8-0=16\\
\bold{Z_{max}=16}"

"\bold{}"